What's the answer and why?
4. For a reaction with a Ho = +50 kJ and a AS。= +100.0 J/K,aGo is the equilibrium constant is and
Answer #1

delta Go = delta HO - T*delta So
= 50,000 J - 298 K * 100 J/K
= 20200 J
= 20.2 KJ

Now use:
delta Go = -R*T*ln Kc
20200 = -8.314*298*ln Kc
Kc = 2.88*10^-4